The product of two consecutive even numbers is 528.
1. Write an equation in words to represent this information.
2. Represent this information algebraically and solve the equation.
3. Find all possible solutions for the two consecutive numbers

Found 2 solutions by rothauserc, jorel1380:
Answer by rothauserc(4718)

1) the product of an even number and the next even number is 528
:
2) let x be the first even number, then
x * (x+2) = 528
x^2 +2x -528 = 0
(x+24) * (x-22) = 0
x = -24 and x = 22
:
3) The two solutions are
22, 24 or
-24, -22
:

Answer by jorel1380(3719)

Let the two consecutive even numbers be n and n+2. Then:
n(n+2)=528
nē+2n-528=0
(n+24)(n-22)=0
n=22 or -24
The pairs of answers are (22,24)(-24,-22)
☺☺☺☺